OpenStudy (anonymous):

If x ÷ -1/6 = 0 , then x = ?

OpenStudy (anonymous):

\[\frac{ x }{- \frac{ 1 }{ 6 } } = 0\] well \[\frac{ x }{- \frac{ 1 }{ 6 } } = -6x\] so -6x = 0 which means x = 0